SO2    +      O3    ------->    SO3    +     O2


EXP         [ A ]        [ B ]       RATE


1           0.25         0.4        0.118


2           0.25         0.20       0.118


3           0.75         0.20       1.062



a. Write down the preliminary equation


b. Explain how you get each of the required orders. EXPLAIN. 


c. Explain, then show how to get the value of "k"  Show math work.


d. Show the complete finished RATE LAW



1
Expert's answer
2021-01-21T03:46:11-0500

"\dfrac{rate2}{rate 1}=""\dfrac{k[SO_2]^m_2[O_3]^n_2}{k[SO_2]^m_1[O_3]^n_1}" }


"\dfrac{0.118}{0.118}=\dfrac{0.25}{0.25}\newline m=1"

SO2 order is 1

"\dfrac{rate3}{rate2}=\dfrac{k[SO_2]^m_3[O_3]^n_3}{k[SO_2]^m_2[O_3]^n_2}"


"\dfrac{1.062}{0.118}=\dfrac{0.75}{0.25}\newline n=3"

O3 order is 3

The value of k is calculated using the following formula